### v1

# Detecting dilute axion stars constrained by fast radio bursts in the solar system via stimulated decay

Abstract
Fast radio bursts (FRBs) can be explained by collapsing axion stars, imposing constraints on the axion parameter space and providing valuable guidance for experimental axion searches. In the traditional post-inflationary model, axion stars could constitute up to $75\%$ of the dark matter component, suggesting that some axion stars may exist within the solar system. Electromagnetic radiation with an angular frequency equal to half the axion mass can stimulate axion decay. Thus, directing a powerful radio beam at an axion star could trigger its stimulated decay, producing a detectable echo. Using this method, we find it is possible to test the existence of dilute axion stars with maximum masses ranging from $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$, as constrained by FRBs, within the solar system. The resulting echo from axion stars constrained by FRBs could be detectable by terrestrial telescopes. Detecting such an echo would confirm the existence of axion stars, unravel the mystery of dark matter, and provide key evidence that some FRBs originate from collapsing axion stars. Furthermore, FRBs produced by axion star collapses could serve as standard candles, aiding in the resolution of the Hubble tension. If no echo is detected using this method, it would place constraints on the abundance of dark matter in the form of dilute axion stars with maximum masses in the range of $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$.
I Introduction
Growing evidence from various observations and theoretical frameworks has established that dark matter constitutes a substantial portion of the Universe’s energy density. Nonetheless, its exact nature remains elusive. The QCD axion Weinberg:1977ma ; Wilczek:1977pj , arising from the Peccei-Quinn mechanism Peccei:1977ur ; Peccei:1977hh developed to resolve the strong CP problem, is a well-motivated candidate for dark matter. Additionally, string theory predicts a spectrum of axion-like particles (ALPs) across a wide mass range, a concept often referred to as the “axiverse” Arvanitaki:2009fg . In this article, we use the term “axions” to encompass both QCD axions and ALPs. Axions can be produced through several mechanisms, such as the misalignment mechanism Preskill:1982cy ; Abbott:1982af ; Dine:1982ah , the decay of string defects Gorghetto:2020qws , and the kinetic misalignment mechanism Co:2019jts . Due to their bosonic nature, axions can achieve extremely high phase space densities, leading to Bose-Einstein condensation (BEC) Sikivie:2009qn and the formation of gravitationally bound structures known as axion stars. For recent reviews, see Refs. Braaten:2019knj ; Visinelli:2021uve ; Zhang:2024bjo .
[…]
Electromagnetic radiation with an angular frequency half the axion mass can stimulate the decay of axions, producing a detectable echo. This phenomenon offers a novel opportunity to search for axion dark matter. By emitting a powerful beam of microwave radiation into space and detecting the resulting echo, axion dark matter may be identified and studied Arza:2019nta ; Arza:2021nec ; Arza:2022dng ; Arza:2023rcs ; Gong:2023ilg . In our previous work, we proposed that the stimulated decay Kephart:1994uy ; Rosa:2017ury ; Caputo:2018vmy ; Di:2023nnb of a dilute axion star associated with planet 9 could be triggered by directing a powerful radio beam at the star, resulting in an echo detectable by terrestrial telescopes Di:2024snm . Such an observation would distinguish a dilute axion star from other planet 9 candidates, such as a PBH or a free-floating planet captured by the solar system Di:2024snm . This method offers a promising avenue for exploring the axion parameter space constrained by FRBs.
This paper is organized as follows. In Sec. II, we briefly introduce the concept of dilute axion stars. In Sec. III, we review the constraints on the axion parameter space imposed by FRBs. Section IV discusses the abundance of dilute axion stars in the Universe. In Sec. V, we analyze the potential for detecting echoes from dilute axion stars, as constrained by FRBs, within the solar system by emitting powerful radio waves into these stars. Finally, Sec. VI presents our conclusions. Throughout this article, we use natural units, where $c=\hbar=1$.
II Dilute Axion Stars
The QCD axion is a hypothetical pseudo-Nambu-Goldstone boson with spin-0, characterized by a small mass $m_{\phi}$, weak self-interactions, and extremely weak couplings to Standard Model particles. The invariance of the axion Lagrangian under the shift symmetry $\phi(x)\rightarrow\phi(x)+2\pi f_{a}$ requires the axion potential $V(\phi)$ to be periodic, such that $V(\phi)=V(\phi+2\pi f_{a})$, where $f_{a}$ is the axion decay constant, representing the energy scale at which the $U(1)_{\rm{PQ}}$ symmetry spontaneously breaks. Performing a series expansion of the axion potential around $\phi=0$, considering only the first two leading terms, the potential can be expressed as
$V(\phi)={1\over 2}m_{\phi}^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4}+{\cal O}\left(\frac{\lambda^{2}\phi^{6}}{6!m_{\phi}^{2}}\right),$ (1)
where $\lambda=-\kappa m_{\phi}^{2}/f_{a}^{2}$ represents the attractive self-interaction coupling constant. The value of $\kappa$ depends on the choice of axion potential. For the instanton potential, $V(\phi)=(m_{\phi}f_{a})^{2}\left[1-\cos\left({\phi}/{f_{a}}\right)\right]$, commonly used in axion phenomenology, $\kappa=1$. For the chiral potential DiVecchia:1980yfw ; GrillidiCortona:2015jxo , $\kappa\simeq 0.34$ GrillidiCortona:2015jxo ; Fujikura:2021omw . Considering the interaction between axions and photons, the general Lagrangian for the axion is given by
${\cal L}$ $=$ ${1\over 2}\partial_{\mu}\phi\partial^{\mu}\phi-{1\over 2}m_{\phi}^{2}\phi^{2}-\frac{\lambda}{4!}\phi^{4}-{\cal O}\left(\frac{\lambda^{2}\phi^{6}}{6!m_{\phi}^{2}}\right)$ (2)
$+\frac{\alpha K}{8\pi f_{a}}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}-{1\over 4}F_{\mu\nu}F^{\mu\nu},$
where $F_{\mu\nu}$ is the electromagnetic field strength tensor, $\tilde{F}^{\mu\nu}$ is its dual, $\alpha$ is the fine-structure constant, and $K$ is a model-dependent constant of order one. Axions are not entirely stable and can decay into two photons via the interaction term ${\cal L}_{int}=\frac{\alpha K}{8\pi f_{a}}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$. This interaction has significant astrophysical implications and provides an experimental window for the search for these elusive particles. The decay rate of an axion at rest into two photons is
$\Gamma_{\phi}$ $=$ $\frac{\alpha^{2}K^{2}m_{\phi}^{3}}{256\pi^{3}f_{a}^{2}}$
$=$ $1.02\times 10^{-50}~{}{\rm{s}^{-1}}K^{2}\left(\frac{m_{\phi}}{10^{-5}~{}\rm{eV}}\right)^{3}\left(\frac{f_{a}}{10^{12}~{}\rm{GeV}}\right)^{-2}.$
For QCD axions, there exists a well-known relationship between the decay constant $f_{a}$ and the axion mass Weinberg:1977ma :
$m_{\phi}$ $=$ $\frac{\sqrt{m_{u}m_{d}}}{m_{u}+m_{d}}\frac{f_{\pi}m_{\pi}}{f_{a}}$ (4)
$\simeq 6\times 10^{-6}~{}{\rm{eV}}\left(\frac{f_{a}}{10^{12}~{}\rm{GeV}}\right)^{-1},$
where $m_{u}\simeq 2.2~{}\rm{MeV}$, $m_{d}\simeq 4.7~{}\rm{MeV}$, and $m_{\pi}\simeq 135~{}\rm{MeV}$ are the masses of the up quark, down quark, and pion, respectively, while $f_{\pi}\simeq 92~{}\rm{MeV}$ is the pion decay constant. This relationship is illustrated in Fig. 1. As bosons, axions can achieve exceptionally high phase space densities, potentially leading to the formation of BECs Sikivie:2009qn . These BECs can result in axion stars, which may manifest in both dilute and dense configurations Chavanis:2017loo ; Visinelli:2017ooc ; Eby:2019ntd . However, dense axion stars may have lifespans too short to significantly impact cosmology as astrophysical objects Braaten:2019knj ; Seidel:1991zh ; Hertzberg:2010yz ; Eby:2015hyx ; Wang:2020zur .
A stable dilute axion star is in equilibrium, where the attractive forces of self-gravity and self-interactions are balanced by the repulsive gradient energy. This equilibrium is maintained as long as the star’s density remains low, thereby minimizing the effects of self-interactions. However, if the mass of the dilute axion star exceeds a critical threshold, determined by the attractive self-interaction coupling constant, this equilibrium becomes disrupted. The maximum mass and corresponding minimum radius of a dilute axion star are given by Chavanis:2011zi ; Chavanis:2011zm
$M_{\rm{max}}\sim 5.073\frac{M_{pl}}{\sqrt{|\lambda|}},~{}~{}~{}~{}~{}~{}R_{\rm{min}}\sim\sqrt{|\lambda|}\frac{M_{pl}}{m_{\phi}}\lambda_{c},$ (5)
where $M_{pl}$ is the Planck mass and $\lambda_{c}$ denotes the Compton wavelength of the axion. When the mass of a dilute axion star exceeds the critical mass given by Eq. (5), either due to merger events Mundim:2010hi ; Cotner:2016aaq ; Schwabe:2016rze ; Eby:2017xaw ; Hertzberg:2020dbk ; Du:2023jxh ; Maseizik:2024qly or through the accretion of axions from the surrounding environment Chen:2020cef ; Chan:2022bkz ; Dmitriev:2023ipv , self-interactions become significant, potentially leading to the collapse of the star Levkov:2016rkk ; Eby:2016cnq ; Fox:2023xgx . Substituting the attractive coupling constant of self-interaction, $\lambda=-\kappa m_{\phi}^{2}/f_{a}^{2}$, and the axion’s Compton wavelength, $\lambda_{c}=2\pi/m_{\phi}$, into Eq. (5), the critical mass and minimum radius are expressed as
$M_{\rm{max}}$ $\sim$ $5.97\times 10^{-12}M_{\odot}~{}\kappa^{-1/2}$ (6)
×$\times\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)^{-1}\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right),$
$R_{\rm{min}}$ $\sim$ $2.41\times 10^{2}~{}{\rm{km}}~{}\kappa^{1/2}$ (7)
×$\times\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)^{-1}\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right)^{-1}.$
These quantities depend not only on the mass and decay constant of the axion but also on the choice of axion potential. For a general dilute axion star, we denote its mass as $M_{\rm{AS}}$ and its radius as $R_{\rm{AS}}$.
III Constraints by fast radio bursts
Since the axion is a pseudoscalar boson, the boson enhancement effect must be considered. The change in photon number density $n_{\lambda^{\prime}}$ within the dilute axion star is governed by the Boltzmann equation Kephart:1994uy :
[…]
×$\times\{f_{\phi}(\textbf{p})[1+f_{\lambda^{\prime}}(\textbf{k}_{1})][1+f_{\lambda^{\prime}}(\textbf{k}_{2})]$
$-f_{\lambda^{\prime}}(\textbf{k}_{1})f_{\lambda^{\prime}}(\textbf{k}_{2})[1+f_{\phi}(\textbf{p})]\},$
where $f_{\phi}$ and $f_{\lambda^{\prime}}$ represent the phase space densities of the axion and photon, respectively. The term $\cal M(\phi\rightarrow\gamma(\lambda^{\prime})\gamma(\lambda^{\prime}))$ refers to the matrix element corresponding to the interaction ${\cal L}_{\text{int}}=\frac{\alpha K}{8\pi f_{a}}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$. The phase space integration is performed using the standard Lorentz-invariant measure, considering the momenta of both the axion and the photons. Assuming the phase space distributions of axions and photons are approximately uniform and isotropic within the dilute axion star, the time evolution of photon number density becomes Kephart:1994uy ; Rosa:2017ury
$\frac{dn_{\gamma}}{dt}=\Gamma_{\phi}\left[2n_{\phi}\left(1+\frac{8\pi^{2}}{m_{\phi}^{3}v}n_{\gamma}\right)-\frac{16\pi^{2}}{3m_{\phi}^{3}}\left(v+\frac{3}{2}\right)n_{\gamma}^{2}\right],$
where $v\sim 1/(2R_{\rm{AS}}m_{\phi})$ represents the maximum axion velocity within the dilute axion star, as estimated using the Heisenberg uncertainty principle.
In Eq. (III), the initial terms within the square brackets represent both spontaneous and stimulated decay. The final term, scaled by a factor of $3/2$, corresponds to the production of “sterile” axions Kephart:1994uy , which can escape from the dilute axion star. Furthermore, photons produced through spontaneous and stimulated decay escape from the dilute axion star at a rate of $\Gamma_{e}=1/{R_{\rm{AS}}}$. By integrating Eq. (III) and accounting for the escaped photons, the coupled differential equations describing the evolution of the number of axions and photons within the dilute axion star are expressed as follows Di:2024tlz
$\frac{dN_{\gamma}}{dt}$ $=$ $2\Gamma_{\phi}N_{\phi}+\frac{12\pi}{m_{\phi}^{3}vR_{\rm{AS}}^{3}}\Gamma_{\phi}N_{\phi}N_{\gamma}$ (10)
$-\frac{2\pi(2v+3)}{m_{\phi}^{3}R_{\rm{AS}}^{3}}\Gamma_{\phi}N_{\gamma}^{2}-\Gamma_{e}N_{\gamma},$
$\frac{dN_{\phi}}{dt}$ $=$ $-\Gamma_{\phi}N_{\phi}-\frac{6\pi}{m_{\phi}^{3}vR_{\rm{AS}}^{3}}\Gamma_{\phi}N_{\phi}N_{\gamma}$ (11)
$+\frac{4\pi v}{m_{\phi}^{3}R_{\rm{AS}}^{3}}\Gamma_{\phi}N_{\gamma}^{2}.$
However, even for dilute axion stars in a critical state, the photon number $N_{\gamma}\simeq(2\Gamma_{\phi}/\Gamma_{e})N_{\phi}$ resulting from spontaneous decay is insufficient to initiate stimulated decay.
When the axion star mass exceeds the critical threshold, it collapses. As the size approaches the critical radius $R_{cr}\simeq 24\pi\Gamma_{\phi}M_{\rm{max}}/m_{\phi}^{3}$ Di:2024tlz , stimulated decay initiates. The luminosity of the collapsing axion star from stimulated decay is given by Di:2024tlz
$L_{\phi}$ $=$ $\frac{1}{2}m_{\phi}N_{\gamma}R_{cr}^{-1}=\frac{N_{\gamma}m_{\phi}^{4}}{48\pi\Gamma_{\phi}M_{\rm{max}}}$
$\sim$ $3.60\times 10^{40}{\rm{erg/s}}\left(\frac{N_{\gamma}}{10^{49}}\right)\left(\frac{m_{\phi}}{10^{-5}{\rm{eV}}}\right)^{2}\left(\frac{f_{a}}{10^{12}{\rm{GeV}}}\right),$
where model-dependent constants $K=1$ and $\kappa=1$ are assumed. Detection of a radio signal consistent with the stimulated decay of a collapsing axion star could serve as a standard candle for constraining the Hubble constant, given the intense luminosity of such a radio burst Di:2024tlz .
Figure 1: The parameter space of axions. The region above the orange solid line is excluded based on constraints from Planck 2018 combined with BAO measurements Caloni:2022uya . Constraints derived from white dwarf observations are represented by the green area Balkin:2022qer , with the latest findings detailed in Ref. Gomez-Banon:2024oux . The purple line corresponds to QCD axions, while the blue region represents the parameter space constrained by FRBs, where the maximum mass of dilute axion stars ranges from $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$.
Interestingly, radio signals from the stimulated decay of collapsing axion stars may already have been detected. FRBs are bright, transient radio emissions lasting only milliseconds Lorimer:2007 ; Keane:2012yh ; Thornton:2013iua , with a frequency range spanning approximately 400 MHz to 8 GHz Petroff:2019tty and total energy outputs typically between $10^{38}$ and $10^{40}$ erg Thornton:2013iua ; Spitler:2014fla . The photon frequency and total energy emitted by a collapsing axion star, for certain parameter ranges, align with those observed in FRBs Di:2024tlz . For constraints on the axion parameter space derived from FRBs, see Ref. Di:2024tlz and Fig. 1. The maximum axion star mass within this constrained parameter space ranges from approximately $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$, as shown in Fig. 1.
IV Abundance of dilute axion stars
The current relic density of axions is described by the equation Di:2023xaw :
$\Omega_{\phi}h^{2}$ $\simeq$ $0.12\left(\frac{g_{\star}(T_{\rm{osc}})}{106.75}\right)^{3/4}\left(\frac{m_{\phi}}{10^{-6}~{}\rm{eV}}\right)^{1/2}$ (13)
×$\times\left(\frac{f_{a}}{5.32\times 10^{12}~{}\rm{GeV}}\right)^{2}.$
This equation, applicable to axions produced via the misalignment mechanism, defines the lower bound of the axion parameter space, as shown in Fig. 1. By combining Eq. (4) with Eq. (13), we obtain the axion mass $m_{\phi}\simeq 1.17\times 10^{-6}$ eV and the decay constant $f_{a}\simeq 5.11\times 10^{12}$ GeV, under the assumption that QCD axions constitute the majority of dark matter. Substituting these values into Eq. (6), the maximum mass of the corresponding dilute axion star is calculated to be approximately $2.61\times 10^{-10}M_{\odot}$. This value coincides with the upper limit of the maximum axion star mass constrained by the parameter space derived from FRBs, as shown in Fig. 1. From Fig. 1, it is evident that the cosmological abundance line intersects the parameter space constrained by FRBs. This overlap significantly enhances the likelihood of axions existing within this region and provides additional support for the misalignment mechanism as a potential source of axions.
In the standard post-inflationary scenario, where $U(1)_{\rm{PQ}}$ symmetry is spontaneously broken after inflation, axion overdensities collapse during the radiation-dominated epoch, forming an early population of miniclusters with masses up to $10^{-12}M_{\odot}$ Hogan:1988mp ; Kolb:1993zz ; Kolb:1995bu . By redshift $z=100$, about $75\%$ of axion dark matter resides within these bound structures Eggemeier:2019khm , with minicluster masses reaching up to $10^{-9}M_{\odot}$ due to mergers. This closely matches the mass range of $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$ for axion stars constrained by FRBs. Dilute axion stars evolve toward a critical state through the accretion of axions from the surrounding environment or via merger events. Once they surpass this critical threshold, they collapse, emitting strong radio signals Di:2023nnb or relativistic axions Levkov:2016rkk . After the collapse, a residual dilute axion star remains, retaining a mass close to its critical value and entering a subcritical state. This suggests that the likelihood of dilute axion stars being in either a critical or subcritical state is relatively high. For simplicity, we will assume that dilute axion stars are in a critical state in the following discussion.
The local dark matter density $\rho_{\rm{DM}}$ near the solar system is estimated to be $0.01M_{\odot}~{}{\rm{pc}}^{-3}\approx 0.4~{}\rm{GeV/cm^{3}}$, as inferred from stellar dynamics on scales greater than approximately 100 parsecs Read:2014qva ; McMillan:2016jtx ; Evans:2018bqy ; deSalas:2020hbh . Using this value, the local number density of dilute axion stars can be calculated as:
$n_{\rm{AS}}$ $=$ $\frac{\Omega_{\rm{AS}}}{\Omega_{\rm{DM}}}\frac{\rho_{\rm{DM}}}{M_{\rm{AS}}}$
$=$ $1.27\times 10^{9}~{}{\rm{pc}}^{-3}\left(\frac{\Omega_{\rm{AS}}}{0.75\Omega_{\rm{DM}}}\right)\left(\frac{6.21\times 10^{-12}M_{\odot}}{M_{\rm{AS}}}\right),$
where $\Omega_{\rm{AS}}/{\Omega_{\rm{DM}}}$ represents the fraction of dilute axion stars within dark matter. Assuming $75\%$ of dark matter is composed of axion stars, a spherical volume with a radius of 1000 AU would contain approximately 605 axion stars with a mass of $6.21\times 10^{-12}M_{\odot}$, or 14 axion stars with a mass of $2.61\times 10^{-10}M_{\odot}$. The number density of dilute axion stars is inversely proportional to their mass: lower-mass axion stars are more numerous, increasing their likelihood of being present in the solar system. This suggests that low-mass dilute axion stars, with maximum masses ranging from $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$, could potentially be detected within the solar system.
V Echo from Dilute Axion Stars
A dilute axion star remains stable because the photon number resulting from spontaneous decay is insufficient to initiate stimulated decay. Now, consider transmitting radio waves with power $P$ to a dilute axion star. The coupled differential equations describing the evolution of the number of axions and photons within the dilute axion star are then modified as follows:
$\frac{dN_{\gamma}}{dt}$ $=$ $2\Gamma_{\phi}N_{\phi}+\frac{12\pi}{m_{\phi}^{3}vR_{\rm{AS}}^{3}}\Gamma_{\phi}N_{\phi}(N_{\gamma}+N_{\gamma 0})$ (15)
$-\frac{2\pi(2v+3)}{m_{\phi}^{3}R_{\rm{AS}}^{3}}\Gamma_{\phi}(N_{\gamma}+N_{\gamma 0})^{2}-\Gamma_{e}N_{\gamma},$
$\frac{dN_{\phi}}{dt}$ $=$ $-\Gamma_{\phi}N_{\phi}-\frac{6\pi}{m_{\phi}^{3}vR_{\rm{AS}}^{3}}\Gamma_{\phi}N_{\phi}(N_{\gamma}+N_{\gamma 0})$ (16)
$+\frac{4\pi v}{m_{\phi}^{3}R_{\rm{AS}}^{3}}\Gamma_{\phi}(N_{\gamma}+N_{\gamma 0})^{2},$
where $N_{\gamma 0}\simeq 2PR_{\rm{AS}}/m_{\phi}$ represents the number of radio wave photons emitted from Earth and transmitted into the dilute axion star. By substituting Eq. (7) with $\kappa=1$ into $N_{\gamma 0}$, we obtain
$N_{\gamma 0}$ $\simeq$ $5.02\times 10^{28}\left(\frac{P}{50~{}\rm{MW}}\right)$ (17)
×$\times\left(\frac{m_{\phi}}{10^{-5}~{}\rm{eV}}\right)^{-2}\left(\frac{f_{a}}{10^{12}~{}\rm{GeV}}\right)^{-1}.$
Stimulated decay dominates over spontaneous decay when the condition $1{6\pi}(N_{\gamma}+N_{\gamma 0})/({m_{\phi}^{3}vR_{\rm{AS}}^{3}})>$ 1 is satisfied. By transmitting radio waves with a power of $P={50~{}\rm{MW}}$ into the dilute axion star, we calculate $N_{\gamma 0}=3.86\times 10^{26}$ and find that ${6\pi}N_{\gamma 0}/({m_{\phi}^{3}vR_{\rm{AS}}^{3}})=2.64\times 10^{15}\gg 1$, using the parameters $m_{\phi}=5\times 10^{-5}$ eV and $f_{a}=5.20\times 10^{12}$ GeV. Thus, radio waves of sufficient power, such as $P=50~{}\rm{MW}$, can effectively induce the stimulated decay of a dilute axion star. The multi-beam relativistic klystron amplifier currently has the capability to generate radio waves with an output power of 1.047 GW sun:2023 , significantly exceeding the required 50 MW. This substantial power output makes it a promising candidate for emitting radio waves to interact with axion stars. In the following discussion, we will use the example of a 50 MW radio wave received by an axion star.
Figure 2: Numerical evolution of the axion, photon, and “sterile” axion numbers within a dilute axion star under a powerful microwave radiation beam of $50{\rm{MW}}$. The solid line represents an axion star with a mass of $M_{\rm{AS}}=2.61\times 10^{-10}M_{\odot}$ and parameters $m_{\phi}=1.17\times 10^{-6}~{}{\rm{eV}}$ and $f_{a}=5.11\times 10^{12}~{}{\rm{GeV}}$. The dashed line corresponds to an axion star with a mass of $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$ and parameters $m_{\phi}=5\times 10^{-5}~{}{\rm{eV}}$ and $f_{a}=5.20\times 10^{12}~{}{\rm{GeV}}$. Timing begins when the axion star is irradiated with radio waves. For $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$, the stimulated decay of the dilute axion star stabilizes after approximately $10^{-4}$ seconds of irradiation, with a photon number of $N_{\gamma}=1.10\times 10^{21}$.
Figure 2 presents two numerical solutions showing the evolution of axion, photon, and “sterile” axion numbers for two representative parameter sets. The solid line corresponds to an axion star with a critical mass of $M_{\rm{AS}}=2.61\times 10^{-10}M_{\odot}$ with parameters $m_{\phi}=1.17\times 10^{-6}$ eV and a decay constant $f_{a}=5.11\times 10^{12}$ GeV, representing the intersection of cosmological abundance and QCD axion constraints in Fig. 1. The dashed line corresponds to an axion star with a critical mass of $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$, with parameters lying within the region constrained by FRBs: $m_{\phi}=5\times 10^{-5}$ eV and $f_{a}=5.20\times 10^{12}$ GeV. For the case $M_{\rm{AS}}=6.21\times 10^{-12}~{}M_{\odot}$, the stimulated decay of the dilute axion star stabilizes after approximately $10^{-4}$ seconds of irradiation, resulting in a photon number $N_{\gamma}=1.10\times 10^{21}$. Considering the photons escaping from the dilute axion star, the luminosity can be expressed as:
$L_{\phi}=\frac{1}{2}m_{\phi}N_{\gamma}\Gamma_{e}.$ (18)
A notable feature of stimulated decay is that, in the axion rest frame, the emitted photons are restricted to propagate either in the same direction as or directly opposite to the incoming photons. However, the velocity distribution of axions within the dilute axion star introduces spatial dispersion of the emitted photons. Consequently, the luminosity flux observed from Earth is roughly given by
$F_{\phi}\sim{L_{\phi}\over 4\pi d^{2}}=\frac{m_{\phi}N_{\gamma}\Gamma_{e}}{8\pi d^{2}},$ (19)
where $d$ represents the distance between the dilute axion star and Earth. For a dilute axion star in its critical state, substituting Eq. (7) with $\kappa=1$ into $\Gamma_{e}=R_{\rm{AS}}^{-1}$, the photon escape rate becomes
$\Gamma_{e}\sim 1.24\times 10^{3}~{}{\rm{s^{-1}}}\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right).$ (20)
By substituting Eq. (20) into Eq. (19), the flux can be expressed as
$F_{\phi}$ $=$ $3.50\times 10^{-35}\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)^{2}\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right)$ (21)
×$\times\left(\frac{N_{\gamma}}{10^{20}}\right)\left(\frac{1000~{}\rm{AU}}{d}\right)^{2}{~{}\rm{W/cm^{2}}}.$
As illustrated in Fig. 1, the critical mass of dilute axion stars constrained by FRBs ranges from approximately $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$. Substituting Eq. (6) with $\kappa=1$ into Eq. (21) for a mass of $6.21\times 10^{-12}M_{\odot}$, the luminosity flux of the dilute axion star is calculated as
$F_{\phi}$ $=$ $3.64\times 10^{-34}\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)^{3}$ (22)
×$\times\left(\frac{N_{\gamma}}{10^{21}}\right)\left(\frac{1000~{}\rm{AU}}{d}\right)^{2}~{}{\rm{W/cm^{2}}},$
corresponding to the blue dot-dashed line with a flux of $F_{\phi}\approx 4.86\times 10^{-32}~{}{\rm{W/cm^{2}}}$ for $d=1000$ AU, as shown in Fig. 3. Similarly, substituting Eq. (6) with $\kappa=1$ into Eq. (21) for a mass of $2.61\times 10^{-10}M_{\odot}$ gives
$F_{\phi}$ $=$ $1.53\times 10^{-32}\left(\frac{m_{\phi}}{10^{-6}~{}{\rm{eV}}}\right)^{3}$ (23)
×$\times\left(\frac{N_{\gamma}}{10^{24}}\right)\left(\frac{1000~{}\rm{AU}}{d}\right)^{2}~{}{\rm{W/cm^{2}}},$
which corresponds to the gray line for $d=1000$ AU, effectively overlapping with the blue dot-dashed line, as shown in Fig. 3. Therefore, the luminosity flux of the echo from dilute axion stars at a distance of 1000 AU, as observed from Earth, is approximately $4.86\times 10^{-32}~{}{\rm{W/cm^{2}}}$, corresponding to an axion star mass range of $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$, as constrained by FRBs. This flux exceeds the sensitivity thresholds of radio telescopes such as SKA, FAST, ngLOBO, and LOFAR with an observation time of 1 hour, as shown in Fig. 3. Thus, the decay signal is detectable by these radio telescopes.
Figure 3: Flux of the dilute axion star at a distance of 1000 AU and the minimum detectable flux of SKA, FAST, ngLOBO, and LOFAR with an observation time of 1 hour. The sensitivity of these telescopes is detailed in Ref. Di:2024snm and its references. Even for LOFAR, the flux of the echoes is one to two orders of magnitude higher than its sensitivity, making the signal easily detectable and distinguishable.
Consequently, the stimulated decay of a dilute axion star within a volume of radius 1000 AU centered around Earth can be triggered by emitting a 50 MW radio beam into the star. The spectral line of the decay signal is nearly monochromatic, with a frequency of $f\simeq m_{\phi}/(4\pi)\approx 1.21(m_{\phi}/10^{-5}{\rm{eV}})~{}{\rm{GHz}}$, making it a distinct characteristic signal. Detection of such an echo would confirm the existence of axion stars, solve the dark matter puzzle, and provide crucial evidence linking some FRBs to axion stars. Furthermore, FRBs originating from collapsing axion stars could serve as standard candles to help resolve the Hubble tension Di:2024tlz . If such echoes are not detected within a 1000 AU volume around Earth using this method, constraints can be placed on the fraction of dark matter composed of dilute axion stars with critical masses constrained by FRBs. This fraction could be roughly less than $5\%$ for a mass of $6.21\times 10^{-12}M_{\odot}$ or $0.1\%$ for a mass of $2.61\times 10^{-10}M_{\odot}$, depending on the axion parameters. This estimate assumes a uniform distribution of axion stars throughout the Universe.
VI Conclusions
Axions are a promising candidate for dark matter. Due to their bosonic nature, axions can achieve very high phase space densities, resulting in BEC and the formation of gravitationally bound structures known as axion stars. Collapsing axion stars with specific parameters may emit millisecond-long radio bursts with peak luminosities of $1.60\times 10^{42}$ erg/s, aligning with the characteristics of observed non-repeating FRBs. If some FRBs are indeed produced by collapsing axion stars, this would impose constraints on the axion parameter space, offering valuable guidance for experimental axion searches. Thus, exploring the axion parameter space in relation to FRBs is crucial, as it relates both to the origin of FRBs and the nature of dark matter.
In the traditional post-inflationary scenario, axion stars could account for up to $75\%$ of dark matter, suggesting that some may exist within the solar system. Building on previous work that explored the possibility of explaining planet 9 as a dilute axion star via stimulated decay induced by a 50 MW radio beam, we propose extending this method to search for dilute axion stars with masses ranging from $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$. These masses are constrained by the characteristics of FRBs, which could potentially be explained by collapsing axion stars. For axions within the parameter space constrained by FRBs, the echo resulting from the stimulated decay of axion stars could be detectable by ground-based radio telescopes. Detection of such an echo would not only confirm the existence of axion stars but also provide key evidence linking axion stars to FRBs. Furthermore, using FRBs from axion star collapses as standard candles could aid in resolving the Hubble tension. If no signal is detected, this method will provide constraints on the abundance of dark matter in the form of dilute axion stars with masses between $6.21\times 10^{-12}M_{\odot}$ and $2.61\times 10^{-10}M_{\odot}$. Under the assumption of a uniform distribution of dilute axion stars across the Universe, the abundance would be limited to less than approximately $5\%$ for masses of $6.21\times 10^{-12}M_{\odot}$ or $0.1\%$ for masses of $2.61\times 10^{-10}M_{\odot}$.
VII Acknowledgments
H. Di would like to thank Lijing Shao for useful discussions. This work was supported by National Natural Science Foundation of China under Grant No. 11947031 and East China University of Technology Research Foundation for Advanced Talents under Grant No. DHBK2019206.

### v4

@@
 In the traditional post-inflationary model, axion stars could constitute up to $75\%$ of the dark matter component, suggesting that some axion stars may exist within the Solar System.
-Electromagnetic radiation with an angular frequency equal to half the axion mass can stimulate axion decay.
+Photons with energy half the axion mass can stimulate axion decay.
 Thus, directing a powerful radio beam at an axion star could trigger its stimulated decay, producing a detectable echo.
@@
 Nonetheless, its exact nature remains elusive.
-The QCD axion Weinberg:1977ma ; Wilczek:1977pj , arising from the Peccei-Quinn mechanism Peccei:1977ur ; Peccei:1977hh developed to resolve the strong CP problem, is a well-motivated candidate for dark matter.
+The QCD axion Weinberg:1977ma ; Wilczek:1977pj , arising from the Peccei-Quinn mechanism Peccei:1977ur ; Peccei:1977hh developed to resolve the strong-$CP$ problem, is a well-motivated candidate for dark matter.
 Additionally, string theory predicts a spectrum of axion-like particles (ALPs) across a wide mass range, a concept often referred to as the “axiverse” Arvanitaki:2009fg .
 In this article, we use the term “axions” to encompass both QCD axions and ALPs.
-Axions can be produced through several mechanisms, such as the misalignment mechanism Preskill:1982cy ; Abbott:1982af ; Dine:1982ah , the decay of string defects Gorghetto:2020qws , and the kinetic misalignment mechanism Co:2019jts .
-Due to their bosonic nature, axions can achieve extremely high phase space densities, leading to Bose-Einstein condensation (BEC) Sikivie:2009qn and the formation of gravitationally bound structures known as axion stars.
-For recent reviews, see Refs.
-Braaten:2019knj ; Visinelli:2021uve ; Zhang:2024bjo .
-[…]
-Electromagnetic radiation with an angular frequency half the axion mass can stimulate the decay of axions, producing a detectable echo.
+Axions can be produced through several mechanisms, such as the misalignment mechanism Preskill:1982cy ; Abbott:1982af ; Dine:1982ah ; Co:2019jts , thermal production Turner:1986tb ; Salvio:2013iaa , the decay of string Davis:1986xc ; Gorghetto:2020qws , and primordial black hole (PBH) evaporation Schiavone:2021imu ; Li:2022xqh ; Mazde:2022sdx .
+Because axions are bosons, they can attain very high phase space densities, which may result in Bose-Einstein condensation (BEC) Sikivie:2009qn and the formation of gravitationally bound objects known as axion stars Braaten:2019knj ; Visinelli:2021uve ; Gorghetto:2024vnp ; Chang:2024fol ; Zhang:2024bjo .
+The unusual orbits of trans-Neptunian objects (TNOs) Brown:2004yy ; Trujillo ; Batygin:2016zsa and the gravitational anomalies detected by the Optical Gravitational Lensing Experiment (OGLE) Mroz:2017mvf remain unexplained.
+One potential explanation is the existence of PBHs Niikura:2019kqi or axion stars Sugiyama:2021xqg with masses in the range of $M\sim 0.5-20M_{\oplus}$, which could account for the OGLE observations.
+To address the TNO orbital anomalies, the planet 9 hypothesis proposes an object with a mass of $5-15M_{\oplus}$ (approximately $1.5-4.5\times 10^{-5}M_{\odot}$) located $300-1000$ AU from the Sun Batygin-PR .
+This hypothetical planet 9 could be explained as a dilute axion star Di:2023xaw that was captured by the Solar System.
+Additionally, collapsing axion stars with specific parameters may emit millisecond-long radio bursts with peak luminosities of approximately $10^{42}$ erg/s, which are consistent with the characteristics of observed non-repeating fast radio bursts (FRBs) Di:2023nnb .
+These collapsing axion stars could serve as novel standard candles Di:2024tlz for constraining the Hubble constant, $H_{0}$, due to their strong, intrinsic luminosity being fixed and dependent solely on the axion mass and decay constant.
+Other studies have also investigated the relationship between axion stars and FRBs Iwazaki:2014wka ; Tkachev:2014dpa ; Raby:2016deh ; Buckley:2020fmh .
+If some FRBs are indeed produced by collapsing axion stars, this would impose constraints on the axion parameter space Di:2024tlz , providing valuable guidance for experimental searches for axions.
+Thus, exploring the axion parameter space in relation to FRBs is crucial, as it offers insight into the origin of FRBs and the nature of dark matter.
+Photons with energy half the axion mass can trigger the stimulated decay Kephart:1994uy ; Rosa:2017ury ; Caputo:2018vmy ; Di:2023nnb ; Dev:2023ijb of axions, producing a echo.
 This phenomenon offers a novel opportunity to search for axion dark matter.
-By emitting a powerful beam of microwave radiation into space and detecting the resulting echo, axion dark matter may be identified and studied Arza:2019nta ; Arza:2021nec ; Arza:2022dng ; Arza:2023rcs ; Gong:2023ilg .
-In our previous work, we proposed that the stimulated decay Kephart:1994uy ; Rosa:2017ury ; Caputo:2018vmy ; Di:2023nnb of a dilute axion star associated with planet 9 could be triggered by directing a powerful radio beam at the star, resulting in an echo detectable by terrestrial telescopes Di:2024snm .
-Such an observation would distinguish a dilute axion star from other planet 9 candidates, such as a PBH or a free-floating planet captured by the solar system Di:2024snm .
+By emitting a radio beam with high power into space and detecting the resulting echo, axion dark matter may be identified and studied Arza:2019nta ; Arza:2021nec ; Arza:2022dng ; Arza:2023rcs ; Gong:2023ilg .
+In our previous work Di:2024snm , we proposed that the stimulated decay of a dilute axion star associated with planet 9 could be triggered by directing a powerful radio beam at the star, resulting in an echo could be detected by current radio telescopes.
+Such an observation would distinguish an axion star from a PBH Scholtz:2019csj or other planet 9 candidates.
 This method offers a promising avenue for exploring the axion parameter space constrained by FRBs.
-This paper is organized as follows.
+This article is organized as follows.
 In Sec.
@@
 In Sec.
-V, we analyze the potential for detecting echoes from dilute axion stars, as constrained by FRBs, within the solar system by emitting powerful radio waves into these stars.
+V, we analyze the potential for detecting echoes from dilute axion stars, as constrained by FRBs, within the Solar System by emitting powerful radio beams into these stars.
 Finally, Sec.
 VI presents our conclusions.
-Throughout this article, we use natural units, where $c=\hbar=1$.
+In this article, we adopt natural units, setting $c=\hbar=1$.
 II Dilute Axion Stars
 The QCD axion is a hypothetical pseudo-Nambu-Goldstone boson with spin-0, characterized by a small mass $m_{\phi}$, weak self-interactions, and extremely weak couplings to Standard Model particles.
-The invariance of the axion Lagrangian under the shift symmetry $\phi(x)\rightarrow\phi(x)+2\pi f_{a}$ requires the axion potential $V(\phi)$ to be periodic, such that $V(\phi)=V(\phi+2\pi f_{a})$, where $f_{a}$ is the axion decay constant, representing the energy scale at which the $U(1)_{\rm{PQ}}$ symmetry spontaneously breaks.
-Performing a series expansion of the axion potential around $\phi=0$, considering only the first two leading terms, the potential can be expressed as
+The invariance of the axion Lagrangian under the shift symmetry $\phi(x)\rightarrow\phi(x)+2\pi f_{a}$ requires the axion potential $V(\phi)$ to be periodic, such that $V(\phi)=V(\phi+2\pi f_{a})$, where $f_{a}$ represents the $U(1)$ symmetry breaking scale, also known as the axion decay constant.
+Performing a series expansion of the axion potential around $\phi=0$, and considering only the first two leading terms, the potential can be expressed as
 $V(\phi)={1\over 2}m_{\phi}^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4}+{\cal O}\left(\frac{\lambda^{2}\phi^{6}}{6!m_{\phi}^{2}}\right),$ (1)
-where $\lambda=-\kappa m_{\phi}^{2}/f_{a}^{2}$ represents the attractive self-interaction coupling constant.
+where $\lambda=-\kappa m_{\phi}^{2}/f_{a}^{2}$ is the coupling constant representing the attractive self-interaction of axions.
 The value of $\kappa$ depends on the choice of axion potential.
@@
 $+\frac{\alpha K}{8\pi f_{a}}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}-{1\over 4}F_{\mu\nu}F^{\mu\nu},$
-where $F_{\mu\nu}$ is the electromagnetic field strength tensor, $\tilde{F}^{\mu\nu}$ is its dual, $\alpha$ is the fine-structure constant, and $K$ is a model-dependent constant of order one.
-Axions are not entirely stable and can decay into two photons via the interaction term ${\cal L}_{int}=\frac{\alpha K}{8\pi f_{a}}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$.
-This interaction has significant astrophysical implications and provides an experimental window for the search for these elusive particles.
-The decay rate of an axion at rest into two photons is
+where $\alpha$ is the fine-structure constant representing the electromagnetic field coupling strength, $K$ is a dimensionless quantity of order one that depends on the axion model, and $F_{\mu\nu}$ and $\tilde{F}^{\mu\nu}$ represent the electromagnetic field tensor and its dual, respectively.
+Due to their coupling with electromagnetic fields, axions can spontaneously decay into photons.
+This interaction has significant astrophysical implications and provides an experimental avenue for searching for these elusive particles.
+The spontaneous decay rate of an axion at rest into two photons is given by
 $\Gamma_{\phi}$ $=$ $\frac{\alpha^{2}K^{2}m_{\phi}^{3}}{256\pi^{3}f_{a}^{2}}$
 $=$ $1.02\times 10^{-50}~{}{\rm{s}^{-1}}K^{2}\left(\frac{m_{\phi}}{10^{-5}~{}\rm{eV}}\right)^{3}\left(\frac{f_{a}}{10^{12}~{}\rm{GeV}}\right)^{-2}.$
+The extremely low decay rate makes it challenging to detect axions through photon signals from their spontaneous decay.
 For QCD axions, there exists a well-known relationship between the decay constant $f_{a}$ and the axion mass Weinberg:1977ma :
@@
 1.
-As bosons, axions can achieve exceptionally high phase space densities, potentially leading to the formation of BECs Sikivie:2009qn .
-These BECs can result in axion stars, which may manifest in both dilute and dense configurations Chavanis:2017loo ; Visinelli:2017ooc ; Eby:2019ntd .
-However, dense axion stars may have lifespans too short to significantly impact cosmology as astrophysical objects Braaten:2019knj ; Seidel:1991zh ; Hertzberg:2010yz ; Eby:2015hyx ; Wang:2020zur .
-A stable dilute axion star is in equilibrium, where the attractive forces of self-gravity and self-interactions are balanced by the repulsive gradient energy.
-This equilibrium is maintained as long as the star’s density remains low, thereby minimizing the effects of self-interactions.
-However, if the mass of the dilute axion star exceeds a critical threshold, determined by the attractive self-interaction coupling constant, this equilibrium becomes disrupted.
-The maximum mass and corresponding minimum radius of a dilute axion star are given by Chavanis:2011zi ; Chavanis:2011zm
+As bosons, axions can achieve exceptionally high phase space densities, potentially leading to the formation of BECs Sikivie:2009qn and the creation of axion stars, which are divided into dilute and dense branches Schiappacasse:2017ham ; Chavanis:2017loo ; Visinelli:2017ooc ; Eby:2019ntd .
+Dilute axion stars are stable against perturbations, while dense axion stars may have lifetimes much shorter than any cosmological timescale Braaten:2019knj ; Seidel:1991zh ; Hertzberg:2010yz ; Eby:2015hyx ; Wang:2020zur .
+For a stable dilute axion star, equilibrium is maintained by the balance between attractive self-gravity and self-interactions, and the repulsive pressure from the Heisenberg uncertainty principle.
+This equilibrium persists as long as the star’s density remains low, minimizing the influence of self-interactions.
+However, once the star’s mass exceeds a critical threshold, determined by the strength of the attractive self-interaction coupling constant, the equilibrium is disrupted.
+The maximum mass $M_{\rm{max}}$ and corresponding minimum radius $R_{\rm{min}}$ of a stable axion star are given by Chavanis:2011zi ; Chavanis:2011zm
 $M_{\rm{max}}\sim 5.073\frac{M_{pl}}{\sqrt{|\lambda|}},~{}~{}~{}~{}~{}~{}R_{\rm{min}}\sim\sqrt{|\lambda|}\frac{M_{pl}}{m_{\phi}}\lambda_{c},$ (5)
-where $M_{pl}$ is the Planck mass and $\lambda_{c}$ denotes the Compton wavelength of the axion.
-When the mass of a dilute axion star exceeds the critical mass given by Eq.
-(5), either due to merger events Mundim:2010hi ; Cotner:2016aaq ; Schwabe:2016rze ; Eby:2017xaw ; Hertzberg:2020dbk ; Du:2023jxh ; Maseizik:2024qly or through the accretion of axions from the surrounding environment Chen:2020cef ; Chan:2022bkz ; Dmitriev:2023ipv , self-interactions become significant, potentially leading to the collapse of the star Levkov:2016rkk ; Eby:2016cnq ; Fox:2023xgx .
+where $\lambda_{c}$ is the Compton wavelength of the axion and $M_{pl}$ is the Planck mass.
+When the mass of an axion star exceeds the critical mass given by Eq.
+(5), either due to merger events Mundim:2010hi ; Cotner:2016aaq ; Schwabe:2016rze ; Eby:2017xaw ; Hertzberg:2020dbk ; Du:2023jxh ; Maseizik:2024qly or through the accretion of axions from the surrounding environment Chen:2020cef ; Chan:2022bkz ; Dmitriev:2023ipv , self-interactions become significant, potentially leading to the collapse of the star Levkov:2016rkk ; Chavanis:2016dab ; Eby:2016cnq ; Fox:2023xgx .
 Substituting the attractive coupling constant of self-interaction, $\lambda=-\kappa m_{\phi}^{2}/f_{a}^{2}$, and the axion’s Compton wavelength, $\lambda_{c}=2\pi/m_{\phi}$, into Eq.
@@
 III Constraints by fast radio bursts
-Since the axion is a pseudoscalar boson, the boson enhancement effect must be considered.
-The change in photon number density $n_{\lambda^{\prime}}$ within the dilute axion star is governed by the Boltzmann equation Kephart:1994uy :
+Despite the extremely low spontaneous decay rate, axion decay rates can be enhanced in the presence of ambient photons, as photons with energy equivalent to half of the axion mass can stimulate decays of axions into two photons.
+The change in photon number density $n_{\lambda^{\prime}}$ within a dilute axion star, resulting from axion decay and inverse decay, is governed by the Boltzmann equation Kephart:1994uy :
 […]
@@
 $-f_{\lambda^{\prime}}(\textbf{k}_{1})f_{\lambda^{\prime}}(\textbf{k}_{2})[1+f_{\phi}(\textbf{p})]\},$
-where $f_{\phi}$ and $f_{\lambda^{\prime}}$ represent the phase space densities of the axion and photon, respectively.
-The term $\cal M(\phi\rightarrow\gamma(\lambda^{\prime})\gamma(\lambda^{\prime}))$ refers to the matrix element corresponding to the interaction ${\cal L}_{\text{int}}=\frac{\alpha K}{8\pi f_{a}}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$.
-The phase space integration is performed using the standard Lorentz-invariant measure, considering the momenta of both the axion and the photons.
-Assuming the phase space distributions of axions and photons are approximately uniform and isotropic within the dilute axion star, the time evolution of photon number density becomes Kephart:1994uy ; Rosa:2017ury
-$\frac{dn_{\gamma}}{dt}=\Gamma_{\phi}\left[2n_{\phi}\left(1+\frac{8\pi^{2}}{m_{\phi}^{3}v}n_{\gamma}\right)-\frac{16\pi^{2}}{3m_{\phi}^{3}}\left(v+\frac{3}{2}\right)n_{\gamma}^{2}\right],$
-where $v\sim 1/(2R_{\rm{AS}}m_{\phi})$ represents the maximum axion velocity within the dilute axion star, as estimated using the Heisenberg uncertainty principle.
+where $\cal M(\phi\rightarrow\gamma(\lambda^{\prime})\gamma(\lambda^{\prime}))$ refers to the matrix element corresponding to the coupling term between axions and photons, $f_{\phi}$ is the phase space density of axions, and $f_{\lambda^{\prime}}$ is the phase space density of photons with helicity $\lambda^{\prime}$.
+After performing the phase space integration for the above equation and considering the two possible helicities of photons, the time evolution of the total photon number density becomes Kephart:1994uy ; Rosa:2017ury
+$\frac{dn_{\gamma}}{dt}=2\Gamma_{\phi}n_{\phi}+\frac{16\pi^{2}}{m_{\phi}^{3}v}\Gamma_{\phi}n_{\phi}n_{\gamma}-\frac{16\pi^{2}}{3m_{\phi}^{3}}\left(v+\frac{3}{2}\right)\Gamma_{\phi}n_{\gamma}^{2},$
+where $v$ is the maximum velocity of an axion within the axion star, which is roughly $1/(2R_{\rm{AS}}m_{\phi})$ as estimated using the uncertainty principle.
+Figure 1: Constraints on the axion parameter space.
+The region above the orange line is excluded based on Planck 2018 and BAO measurements Caloni:2022uya .
+A recent study Springmann:2024ret places a model-independent bound on QCD axions from supernova 1987A, which is stronger than the bounds from Planck and BAO.
+Constraints from white dwarf observations are represented by the green region Balkin:2022qer , with recent updates provided in Ref.
+Gomez-Banon:2024oux .
+The purple line indicates constraints for QCD axions, while the blue region corresponds to the parameter space constrained by FRBs, where the maximum mass of dilute axion stars ranges from $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$.
 In Eq.
-(III), the initial terms within the square brackets represent both spontaneous and stimulated decay.
-The final term, scaled by a factor of $3/2$, corresponds to the production of “sterile” axions Kephart:1994uy , which can escape from the dilute axion star.
-Furthermore, photons produced through spontaneous and stimulated decay escape from the dilute axion star at a rate of $\Gamma_{e}=1/{R_{\rm{AS}}}$.
-By integrating Eq.
-(III) and accounting for the escaped photons, the coupled differential equations describing the evolution of the number of axions and photons within the dilute axion star are expressed as follows Di:2024tlz
+(III), the first term represents spontaneous axion decay, while the second term accounts for the stimulated decay of axions in the presence of ambient photons.
+The third term corresponds to the inverse decay process, in which photons are converted back into axions.
+The component with a factor of $3/2$ corresponds to the generation of “sterile” axions with high velocity that escape from the axion star Kephart:1994uy .
+Additionally, photons produced through spontaneous axion decay and stimulated decay in the presence of ambient photons can also escape from the axion star, with an escape rate of $\Gamma_{e}=1/{R_{\rm{AS}}}$.
+Integrating Eq.
+(III) and considering the photons that escape from the axion star, the coupled differential equations describing the evolution of the number of axions and photons within the axion star are given by Di:2024tlz
 $\frac{dN_{\gamma}}{dt}$ $=$ $2\Gamma_{\phi}N_{\phi}+\frac{12\pi}{m_{\phi}^{3}vR_{\rm{AS}}^{3}}\Gamma_{\phi}N_{\phi}N_{\gamma}$ (10)
@@
 Detection of a radio signal consistent with the stimulated decay of a collapsing axion star could serve as a standard candle for constraining the Hubble constant, given the intense luminosity of such a radio burst Di:2024tlz .
-Figure 1: The parameter space of axions.
-The region above the orange solid line is excluded based on constraints from Planck 2018 combined with BAO measurements Caloni:2022uya .
-Constraints derived from white dwarf observations are represented by the green area Balkin:2022qer , with the latest findings detailed in Ref.
-Gomez-Banon:2024oux .
-The purple line corresponds to QCD axions, while the blue region represents the parameter space constrained by FRBs, where the maximum mass of dilute axion stars ranges from $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$.
 Interestingly, radio signals from the stimulated decay of collapsing axion stars may already have been detected.
@@
 1.
-The maximum axion star mass within this constrained parameter space ranges from approximately $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$, as shown in Fig.
+The constraints on the axion parameter space provide valuable guidance for experimental searches for axions.
+The maximum axion star mass within this constrained parameter space, assuming $\kappa=1$, ranges from approximately $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$, as shown in Fig.
 1.
@@
 (4) with Eq.
-(13), we obtain the axion mass $m_{\phi}\simeq 1.17\times 10^{-6}$ eV and the decay constant $f_{a}\simeq 5.11\times 10^{12}$ GeV, under the assumption that QCD axions constitute the majority of dark matter.
+(13), we obtain the axion mass $m_{\phi}\simeq 1.17\times 10^{-6}$ eV and the decay constant $f_{a}\simeq 5.11\times 10^{12}$ GeV, assuming that QCD axions constitute the majority of dark matter.
 Substituting these values into Eq.
-(6), the maximum mass of the corresponding dilute axion star is calculated to be approximately $2.61\times 10^{-10}M_{\odot}$.
-This value coincides with the upper limit of the maximum axion star mass constrained by the parameter space derived from FRBs, as shown in Fig.
-1.
-From Fig.
-1, it is evident that the cosmological abundance line intersects the parameter space constrained by FRBs.
-This overlap significantly enhances the likelihood of axions existing within this region and provides additional support for the misalignment mechanism as a potential source of axions.
+(6) and taking $\kappa=1$, we calculate the maximum mass of the corresponding dilute axion star to be approximately $2.61\times 10^{-10}M_{\odot}$.
+This value coincides with the upper limit of the maximum axion star mass, as constrained by the parameter space derived from FRBs.
+As shown in Fig.
+1, the cosmological abundance line intersects this parameter space.
+This overlap significantly increases the likelihood of axions existing in this region and further supports the misalignment mechanism as a potential source of axions.
 In the standard post-inflationary scenario, where $U(1)_{\rm{PQ}}$ symmetry is spontaneously broken after inflation, axion overdensities collapse during the radiation-dominated epoch, forming an early population of miniclusters with masses up to $10^{-12}M_{\odot}$ Hogan:1988mp ; Kolb:1993zz ; Kolb:1995bu .
 By redshift $z=100$, about $75\%$ of axion dark matter resides within these bound structures Eggemeier:2019khm , with minicluster masses reaching up to $10^{-9}M_{\odot}$ due to mergers.
-This closely matches the mass range of $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$ for axion stars constrained by FRBs.
+This closely matches the mass range of $6.21\times 10^{-12}M_{\odot}$ to $2.61\times 10^{-10}M_{\odot}$ for critical axion stars constrained by FRBs.
 Dilute axion stars evolve toward a critical state through the accretion of axions from the surrounding environment or via merger events.
@@
 A dilute axion star remains stable because the photon number resulting from spontaneous decay is insufficient to initiate stimulated decay.
-Now, consider transmitting radio waves with power $P$ to a dilute axion star.
-The coupled differential equations describing the evolution of the number of axions and photons within the dilute axion star are then modified as follows:
+However, by transmitting a powerful radio beam with power $P$ to the axion star, with an angular frequency of photons equal to half the axion mass, the stimulated decay of axions can be significantly enhanced.
+As a result, the coupled differential equations given by Eq.
+(10) and Eq.
+(11) are modified as follows:
 $\frac{dN_{\gamma}}{dt}$ $=$ $2\Gamma_{\phi}N_{\phi}+\frac{12\pi}{m_{\phi}^{3}vR_{\rm{AS}}^{3}}\Gamma_{\phi}N_{\phi}(N_{\gamma}+N_{\gamma 0})$ (15)
@@
 $+\frac{4\pi v}{m_{\phi}^{3}R_{\rm{AS}}^{3}}\Gamma_{\phi}(N_{\gamma}+N_{\gamma 0})^{2},$
-where $N_{\gamma 0}\simeq 2PR_{\rm{AS}}/m_{\phi}$ represents the number of radio wave photons emitted from Earth and transmitted into the dilute axion star.
-By substituting Eq.
-(7) with $\kappa=1$ into $N_{\gamma 0}$, we obtain
-$N_{\gamma 0}$ $\simeq$ $5.02\times 10^{28}\left(\frac{P}{50~{}\rm{MW}}\right)$ (17)
+where $N_{\gamma 0}\simeq 2PR_{\rm{AS}}/m_{\phi}$ represents the number of photons from the radio beam transmitted into the axion star. By substituting Eq. (7) into $N_{\gamma 0}$, we obtain
+$N_{\gamma 0}$ $\simeq$ $5.02\times 10^{28}~{}\kappa^{1/2}\left(\frac{P}{50~{}\rm{MW}}\right)$ (17)
 ×$\times\left(\frac{m_{\phi}}{10^{-5}~{}\rm{eV}}\right)^{-2}\left(\frac{f_{a}}{10^{12}~{}\rm{GeV}}\right)^{-1}.$
 Stimulated decay dominates over spontaneous decay when the condition $1{6\pi}(N_{\gamma}+N_{\gamma 0})/({m_{\phi}^{3}vR_{\rm{AS}}^{3}})>$ 1 is satisfied.
-By transmitting radio waves with a power of $P={50~{}\rm{MW}}$ into the dilute axion star, we calculate $N_{\gamma 0}=3.86\times 10^{26}$ and find that ${6\pi}N_{\gamma 0}/({m_{\phi}^{3}vR_{\rm{AS}}^{3}})=2.64\times 10^{15}\gg 1$, using the parameters $m_{\phi}=5\times 10^{-5}$ eV and $f_{a}=5.20\times 10^{12}$ GeV.
+By transmitting a powerful radio beam with a power of $50~{}\rm{MW}$ into the axion star, we obtain $N_{\gamma 0}=3.86\times 10^{26}$ and find that ${6\pi}N_{\gamma 0}/({m_{\phi}^{3}vR_{\rm{AS}}^{3}})=2.64\times 10^{15}\gg 1$, using the parameters $m_{\phi}=5\times 10^{-5}$ eV, $f_{a}=5.20\times 10^{12}$ GeV and taking the model-dependent constant $\kappa=1$.
 Thus, radio waves of sufficient power, such as $P=50~{}\rm{MW}$, can effectively induce the stimulated decay of a dilute axion star.
-The multi-beam relativistic klystron amplifier currently has the capability to generate radio waves with an output power of 1.047 GW sun:2023 , significantly exceeding the required 50 MW.
+The multi-beam relativistic klystron amplifier currently has the capability to output radio waves with a power of 1.047 GW sun:2023 , significantly exceeding the required 50 MW.
 This substantial power output makes it a promising candidate for emitting radio waves to interact with axion stars.
 In the following discussion, we will use the example of a 50 MW radio wave received by an axion star.
-Figure 2: Numerical evolution of the axion, photon, and “sterile” axion numbers within a dilute axion star under a powerful microwave radiation beam of $50{\rm{MW}}$.
+Figure 2: Evolving axion, photon, and “sterile” axion numbers inside a dilute axion star under a $50{\rm{MW}}$ radio beam.
 The solid line represents an axion star with a mass of $M_{\rm{AS}}=2.61\times 10^{-10}M_{\odot}$ and parameters $m_{\phi}=1.17\times 10^{-6}~{}{\rm{eV}}$ and $f_{a}=5.11\times 10^{12}~{}{\rm{GeV}}$.
 The dashed line corresponds to an axion star with a mass of $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$ and parameters $m_{\phi}=5\times 10^{-5}~{}{\rm{eV}}$ and $f_{a}=5.20\times 10^{12}~{}{\rm{GeV}}$.
-Timing begins when the axion star is irradiated with radio waves.
+Timing begins when the axion star is irradiated with radio beam.
 For $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$, the stimulated decay of the dilute axion star stabilizes after approximately $10^{-4}$ seconds of irradiation, with a photon number of $N_{\gamma}=1.10\times 10^{21}$.
-Figure 2 presents two numerical solutions showing the evolution of axion, photon, and “sterile” axion numbers for two representative parameter sets.
-The solid line corresponds to an axion star with a critical mass of $M_{\rm{AS}}=2.61\times 10^{-10}M_{\odot}$ with parameters $m_{\phi}=1.17\times 10^{-6}$ eV and a decay constant $f_{a}=5.11\times 10^{12}$ GeV, representing the intersection of cosmological abundance and QCD axion constraints in Fig.
+Figure 2 presents two numerical solutions illustrating the evolution of the axion, photon, and “sterile” axion numbers for two representative parameter sets, assuming the model-dependent constants $K=1$ and $\kappa=1$.
+The solid line represents an axion star with a critical mass of $M_{\rm{AS}}=2.61\times 10^{-10}M_{\odot}$ with parameters $m_{\phi}=1.17\times 10^{-6}$ eV and a decay constant $f_{a}=5.11\times 10^{12}$ GeV, representing the intersection of cosmological abundance and QCD axion constraints in Fig.
 1.
-The dashed line corresponds to an axion star with a critical mass of $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$, with parameters lying within the region constrained by FRBs: $m_{\phi}=5\times 10^{-5}$ eV and $f_{a}=5.20\times 10^{12}$ GeV.
+The dashed line represents an axion star with a critical mass of $M_{\rm{AS}}=6.21\times 10^{-12}M_{\odot}$, with parameters lying within the region constrained by FRBs: $m_{\phi}=5\times 10^{-5}$ eV and $f_{a}=5.20\times 10^{12}$ GeV.
 For the case $M_{\rm{AS}}=6.21\times 10^{-12}~{}M_{\odot}$, the stimulated decay of the dilute axion star stabilizes after approximately $10^{-4}$ seconds of irradiation, resulting in a photon number $N_{\gamma}=1.10\times 10^{21}$.
@@
 $L_{\phi}=\frac{1}{2}m_{\phi}N_{\gamma}\Gamma_{e}.$ (18)
-A notable feature of stimulated decay is that, in the axion rest frame, the emitted photons are restricted to propagate either in the same direction as or directly opposite to the incoming photons.
-However, the velocity distribution of axions within the dilute axion star introduces spatial dispersion of the emitted photons.
+A notable feature of axion stimulated decay is that, in the axion rest frame, the emitted photons are restricted to propagate either in the same direction as or directly opposite to the incoming photons.
+However, within the dilute axion star, the velocity distribution of the axions leads to spatial dispersion of the emitted photons.
 Consequently, the luminosity flux observed from Earth is roughly given by
 $F_{\phi}\sim{L_{\phi}\over 4\pi d^{2}}=\frac{m_{\phi}N_{\gamma}\Gamma_{e}}{8\pi d^{2}},$ (19)
-where $d$ represents the distance between the dilute axion star and Earth.
+where $d$ denotes the distance between the dilute axion star and Earth.
 For a dilute axion star in its critical state, substituting Eq.
-(7) with $\kappa=1$ into $\Gamma_{e}=R_{\rm{AS}}^{-1}$, the photon escape rate becomes
-$\Gamma_{e}\sim 1.24\times 10^{3}~{}{\rm{s^{-1}}}\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right).$ (20)
-By substituting Eq. (20) into Eq. (19), the flux can be expressed as
-$F_{\phi}$ $=$ $3.50\times 10^{-35}\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)^{2}\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right)$ (21)
+(7) into $\Gamma_{e}=R_{\rm{AS}}^{-1}$, the photon escape rate becomes
+$\Gamma_{e}\sim 1.24\times 10^{3}~{}{\rm{s^{-1}}}~{}\kappa^{-1/2}\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right).$
+By substituting Eq. (V) into Eq. (19), the flux can be expressed as
+$F_{\phi}$ $=$ $3.50\times 10^{-35}~{}\kappa^{-1/2}\left(\frac{m_{\phi}}{10^{-5}~{}{\rm{eV}}}\right)^{2}\left(\frac{f_{a}}{10^{12}~{}{\rm{GeV}}}\right)$ (21)
 ×$\times\left(\frac{N_{\gamma}}{10^{20}}\right)\left(\frac{1000~{}\rm{AU}}{d}\right)^{2}{~{}\rm{W/cm^{2}}}.$
@@
 Substituting Eq.
-(6) with $\kappa=1$ into Eq.
+(6) into Eq.
 (21) for a mass of $6.21\times 10^{-12}M_{\odot}$, the luminosity flux of the dilute axion star is calculated as
@@
 Similarly, substituting Eq.
-(6) with $\kappa=1$ into Eq.
+(6) into Eq.
 (21) for a mass of $2.61\times 10^{-10}M_{\odot}$ gives
@@
 Thus, the decay signal is detectable by these radio telescopes.
-Figure 3: Flux of the dilute axion star at a distance of 1000 AU and the minimum detectable flux of SKA, FAST, ngLOBO, and LOFAR with an observation time of 1 hour.
+Figure 3: Axion star flux at 1000 AU and the sensitivity of SKA, FAST, ngLOBO, and LOFAR for 1 hour observations.
 The sensitivity of these telescopes is detailed in Ref.
@@
 Consequently, the stimulated decay of a dilute axion star within a volume of radius 1000 AU centered around Earth can be triggered by emitting a 50 MW radio beam into the star.
-The spectral line of the decay signal is nearly monochromatic, with a frequency of $f\simeq m_{\phi}/(4\pi)\approx 1.21(m_{\phi}/10^{-5}{\rm{eV}})~{}{\rm{GHz}}$, making it a distinct characteristic signal.
+The decay signal produces a spectral line that is almost monochromatic, with a frequency given by $f\simeq m_{\phi}/(4\pi)\approx 1.21(m_{\phi}/10^{-5}{\rm{eV}})~{}{\rm{GHz}}$, which serves as a distinctive feature of the signal.
 Detection of such an echo would confirm the existence of axion stars, solve the dark matter puzzle, and provide crucial evidence linking some FRBs to axion stars.
@@
 Axions are a promising candidate for dark matter.
-Due to their bosonic nature, axions can achieve very high phase space densities, resulting in BEC and the formation of gravitationally bound structures known as axion stars.
-Collapsing axion stars with specific parameters may emit millisecond-long radio bursts with peak luminosities of $1.60\times 10^{42}$ erg/s, aligning with the characteristics of observed non-repeating FRBs.
+Due to their bosonic properties, they can reach exceptionally high phase space densities, which may lead to BEC and the formation of gravitationally bound objects known as axion stars.
+Collapsing axion stars with specific parameters may emit millisecond-long radio bursts with peak luminosities of approximately $10^{42}$ erg/s, which are consistent with the characteristics of observed non-repeating FRBs.
 If some FRBs are indeed produced by collapsing axion stars, this would impose constraints on the axion parameter space, offering valuable guidance for experimental axion searches.
@@
 Under the assumption of a uniform distribution of dilute axion stars across the Universe, the abundance would be limited to less than approximately $5\%$ for masses of $6.21\times 10^{-12}M_{\odot}$ or $0.1\%$ for masses of $2.61\times 10^{-10}M_{\odot}$.
-VII Acknowledgments
+Acknowledgements.
 H.
-Di would like to thank Lijing Shao for useful discussions.
+D.
+would like to thank Lijing Shao for useful discussions.
 This work was supported by National Natural Science Foundation of China under Grant No.